OpenStudy (anonymous):

11 Kariharo 3rd grader-Explain very basic please. This is the question- a clock starts at 5:05. What time is it after the minute hand rotates 12 degrees?

OpenStudy (anonymous):

There are 360 degrees in a circle, and a clock is a circle, right? So, divide 12 by 360. You get 1 out of 30. If the minute hand rotates 12 degrees, it moves two minutes, since 30 times 2 = 60, and there are 60 minutes in a hour, which means that 360 degrees on a clock is 60 minutes. Sorry, does that make sense?

OpenStudy (triciaal):

a clock is a circle 360° one revolution is 60 minutes 12° out of 360° = same fraction as the number of minutes out of 60 let the number of minutes be x 12/360 = x/60 x = 12*60/360 simplify by dividing by common factors x = 2 minutes start at 5:05 and 2 minutes later means 5:05 + 0:02 = ?